# Winding quotients for virtual period maps of rank 1

### v4

@@ Abstract.
We illustrate a rank 1 model of virtual period maps and their associated winding quotients. The winding quotient is a new notion that appears in the recent study of virtual period maps. It requires a reformulation of the classical inversion problem for the classical period maps. We answer to the new inversion problem by introducing a $q$-multiplicatively periodic function, whose pull-back to the winding covering space is the Weierstrass $\wp$-function up to a correction by Eisenstein series $E_{2}$. This function appears also in the study of mathematical physics as the propagator on elliptic curves.
@@ 1. Introduction
The classical theory of elliptic integral of the first kind for a Weierstrass family of elliptic curves $X\to S$ (e.g. [Si] Chap.1) defines a period map, which is a multi-valent map from the complement of the discriminant divisor $S\setminus D$ of the base space $S$ of the family to its range domain (so called, period domain). The range is isomorphic to the monodromy covering of $S\setminus D$, i.e. the quotient space of the universal covering $(S\setminus D)^{\sim}$ divided by the kernel of the monodromy representation of $\pi_{1}(S\setminus D,*)$. Then, the inversion problem for the period map ([ibid] 1.13.) is to ask for a description of the inverse map from the range to the defining domain $S\setminus D$, and its answer were given by Eisenstein series (see [ibid] Section13)
The theory of period integrals for primitive forms [S1] has generalized those classical theory of period integrals to wider classes of geometric families $X\to S$. In particular, the elliptic integrals of Weierstrass type, Jacobi type and Hesse type are explained by the primitive forms associated to the root systems of types $\mathrm{A}_{2},\mathrm{B}_{2}$ and $\mathrm{G}_{2}$, respectively ([S2])). Recently, we find that even for hyperbolic root systems of rank 2, which do not have geometric origin, admit primitive forms and associated period maps ([S3]). We call such period map “virtual” since it does not have corresponding geometric family $X\to S$. Then, in these virtual period maps, we observe some new phenomenon. Namely, an infinite cyclic group, called the winding group, acts on the defining domain and on the monodromy covering space of the period map. It seems to ask for a reformulation of the classical inversion problem.
The purpose of the present note is to illustrate such winding phenomenon by an elementary model of virtual period maps of rank 1. We observe that the winding group, acting on the monodromy covering space, acts also on the defining domain of the period map so that we obtain the winded quotient of the defining domain. Then, the classical inversion problem is reformulated in this new setting to a problem to describe the morphism from the winding quotient of the range to the winding quotient of the domain. We solve the new problem by using a $q$-multiplicatively periodic function $\mathcal{W}$ on the range whose pull-back to the monodromy covering space (i.e. the winding covering space of the range) is the Fourier expansion part of Weierstrass $\wp$-function up to a correction term by Eisenstein series $E_{2}$ studied by Borcherds [B].
@@
Actually, the pull-back of the function $\mathcal{W}$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves and the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [LZ] 3.3), or as the genus-one Bergman kernel function $B(z_{1},z_{2})$ in the topological recursion by Eynard [E] p.339 Examples (c.f. also Tyurin [T] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S3, GS].
@@ 3. Lifting of the map $z^{\tau}$ to the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$
The naive “invertion problem” in this setting is to give a map from $\mathbb{C}_{w}^{\times}$ to $\mathbb{C}_{z}^{\times}$ making the diagram commutative. However, there does not exists a univalent holomorphic map which makes the diagram commutative, since the analytic continuation of any local inversion map becomes automatically the multi-valued function $z=w^{1/\tau}$ on $\mathbb{C}_{w}^{\times}$. That is, a naive solution of the classical inversion problem for the virtual period map $z^{\tau}$ by a univalent map from the range to the domain does not exist.
@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
where the scaling factor $\frac{\tau}{2\pi\sqrt{-1}}$ comes from the scaling of the group $\Gamma_{g}+\Gamma_{w}=\frac{2\pi\sqrt{-1}}{\tau}(\mathbb{Z}+\mathbb{Z}\tau)$ relative to the lattice $\mathbb{Z}+\mathbb{Z}\tau$ for the $\wp$-function. Thus, the new inversion problem asks to describe the pair coordinates $(\wp(u,\tau),\wp^{\prime}(u,\tau))$ of the point $e\in E_{\tau}$ in terms of $w=\exp(\tau\mathrm{Log}z)=\exp(2\pi\sqrt{-1}u)$ and $\tau$.  [^3] The realization of $E_{\tau}$ as a plane curve is not unique, but there are severals (e.g. [S2]), and in the present note, we have chosen just one of them, i.e. $\mathrm{A}_{2}$-type realization. It would be also interesting to study the other cases $\mathrm{B}_{2}$-type and $\mathrm{G}_{2}$-type.
6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
In order to answer to the new inversion problem, in the present section, we study a meromorphic function $\mathcal{W}$ on $\mathbb{C}_{w}^{\times}$ which is invariant by the multiplicative action of $q^{\mathbb{Z}}:=\exp(\tau\Gamma_{g})=\exp(2\pi\sqrt{-1}\tau\mathbb{Z})$ on $\mathbb{C}_{w}^{\times}$.
Following the classical idea of the construction of the $\wp$-function by its partial fractional expansion on the $u$-plane $\mathbb{C}_{u}$ (which is additively doubly periodic, having second order poles at $\mathbb{Z}+\mathbb{Z}\tau$), we construct a meromorphic function $\mathcal{W}$ by its partial fractional expansion on $w\in\mathbb{C}_{w}^{\times}$ which is multiplicatively periodic by the periods $q^{\mathbb{Z}}$, i.e. $\mathcal{W}(wq,q)=\mathcal{W}(w,q)$, having second order poles on the set $\cup_{m\in\mathbb{Z}}\{1-wq^{m}=0\}$.
Let us consider an infinite (formal) sum of rational functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$:
$(1)\qquad\qquad\qquad\qquad\qquad\qquad\mathcal{W}(w,q)\ \ :=\ \ \sum_{m\in\mathbb{Z}}\frac{wq^{m}}{(1-wq^{m})^{2}}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad$
where the $m$-th term $\frac{wq^{m}}{(1-wq^{m})^{2}}$ ($m\in\mathbb{Z}$) is a rational function having the second order pole at $w=q^{-m}$ which is invariant by the involution: $(w,q)\mapsto(w^{-1},q^{-1})$.
Assertion A. The sum obtained by removing the $m$-th term from $(1)$ converges uniformly on any compact subset of the domain $\mathcal{U}_{m}:=\{(w,q)\in\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}\mid|q|^{-m+1}<|w|<|q|^{-m-1}\}$. So the sum $(1)$ defines a meromorphic function on $\mathcal{U}_{m}$.
Since we have the open covering $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}=\cup_{m\in\mathbb{Z}}\ \mathcal{U}_{m}$, the sum $(1)$ converges to a meromorphic function on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, which we shall denote again by $\mathcal{W}(w,q)$.
The function $\mathcal{W}(w,q)$ is $q$-multiplicatively periodic.
$(2)\qquad\qquad\qquad\quad\qquad\qquad\qquad\mathcal{W}(wq,q)\ \ =\ \ \mathcal{W}(w,q)\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
@@ 6. $q$-multiplicatively periodic function $\mathcal{W}(w,q)$
We first decompose the sum in two parts:
$(3)\qquad\qquad\mathcal{W}(w,q)-\frac{wq^{m}}{(1-wq^{m})^{2}}\ =\ \ \sum_{n\in\mathbb{Z}_{>m}}\frac{wq^{n}}{(1-wq^{n})^{2}}+\sum_{n\in\mathbb{Z}_{>-m}}\frac{w^{-1}q^{n}}{(1-w^{-1}q^{n})^{2}}\qquad\qquad\qquad\qquad\qquad$
where the second part is: $\sum_{n\in\mathbb{Z}_{<m}}\frac{wq^{n}}{(1-wq^{n})^{2}}=\sum_{n\in\mathbb{Z}_{>-m}}\frac{w^{-1}q^{n}}{(1-w^{-1}q^{n})^{2}}$ (the involutivity).
Notation. For any compact subset $K$ of the domain $\mathcal{U}_{m}$ and for any complex valued continuous function $f$ on $\mathcal{U}_{m}$, we set $f|\!|_{K}:=\mathrm{sup}\{|f(w,q)|\mid(w,q)\in K\}$.
We next show that, for the given $m\in\mathbb{Z}$ and $K$, there exists a positive constants $C\in\mathbb{R}_{>0}$ such that
@@
$\qquad C:=\mathrm{sup}\big\{2,\ |\!|\frac{1}{1-wq^{n}}|\!|_{K}\ \text{for}\ m<n<N,\ |\!|\frac{1}{1-w^{-1}q^{n}}|\!|_{K}\ \text{for}\ -m<n<N\ \big\}\quad\Box\!\!$
Thus, we finally have the following majorations for each parts of $(3)$
[…]
which implies the compact uniform convergence of the sum in the right hand side of $(3)$. So, the sum gives a holomorphic function on the domain $\mathcal{U}_{m}$. Thus the sum (1) defines a meromorphic function on $\mathcal{U}_{m}$, having the pole of order two along $1\!-\!wq^{m}\!=\!0$.
The identification of the functions on $\mathcal{U}_{m}$ and on $\mathcal{U}_{m+1}$ on the intersection $\mathcal{U}_{m}\cap\mathcal{U}_{m+1}$ is achieved by separating two terms $\frac{wq^{m}}{(1-wq^{m})^{2}}$ and $\frac{wq^{m+1}}{(1-wq^{m+1})^{2}}$ from each sum so that the remaining sums are the same. Thus the global function $\mathcal{W}(w,q)$ is defined.
The function $\mathcal{W}(wq,q)$ is defined by one shift of the summation index of the sum $(1)$. The compact uniform convergence of the sum (1) implies that we can change the summation order, implying the equality (2) as meromorphic functions. ∎
Assertion B. The function $\mathcal{W}$ on $\mathbb{C}_{w}^{\times}$ and the $\wp$-function on $\mathbb{C}_{\mathrm{Log}z}$ are related by the following pull-back relations.
$\begin{array}[]{ccl}\wp(u,\tau)&=&(2\pi\sqrt{-1})^{2}\ \mathcal{W}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)\\ \wp^{\prime}(u,\tau)&=&(2\pi\sqrt{-1})^{3}\ \mathcal{W}^{\prime}(w,q)\end{array}$
where we set $w=\exp(2\pi\sqrt{-1}u),\ q=\exp(2\pi\sqrt{-1}\tau)$, and $E_{2}(\tau)=\frac{1}{\pi^{2}}\sum_{m}\big(\sum_{n}^{\prime}\frac{1}{(m\tau+n)^{2}}\big)$ is the Eisenstein series of weight 2 (here $\sum_{n}^{\prime}$ means n=0n\!=\!0 is omitted if m=0m\!=\!0, see Zagier [Z] p. 19).
@@
For each fixed $\tau\in\mathbb{H}$, the substitution of $w=\exp(2\pi\sqrt{-1}u)$ and $q=\exp(2\pi\sqrt{-1}\tau)$ to $\mathcal{W}(w,q)$ is a doubly periodic meromorphic function in $u$ having poles at the lattice $\mathbb{Z}+\mathbb{Z}\tau$, and the principal part of its Laurent expansion at $u=0$ is equal to $\frac{1}{(2\pi\sqrt{-1})^{2}}\frac{1}{u^{2}}$. Then, the difference of the $\wp(u,\tau)$ and the substitution $(2\pi\sqrt{-1})^{2}\mathcal{W}(\exp(2\pi\sqrt{-1}u),\exp(2\pi\sqrt{-1}\tau))$ is a doubly periodic function on $u\in\mathbb{C}_{u}$ without poles so that it is a constant function on $u$. That is, the difference is a function depending only on $\tau$.
@@
That is, the pull-back relation holds on the open set $\mathbb{C}_{u}\times\mathbb{H}_{\tau}\setminus\cup_{m\in\mathbb{Z}}\{\Im(u+m\tau)=0\}$. Clearly, this set projects surjectively onto $\mathbb{H}_{\tau}$. ∎

### v5

@@ Abstract.
We illustrate a rank 1 model of virtual period maps and their associated winding quotients. The winding quotient is a new notion that appears in the recent study of virtual period maps. It requires a reformulation of the classical inversion problem for the classical period maps. We answer to the new inversion problem by introducing a $q$-multiplicatively periodic function, whose pull-back to the winding covering space is the Weierstrass $\wp$-function up to a correction by Eisenstein series $E_{2}$. This function appears as the propagator on elliptic curves in mathematical physics.
@@ 1. Introduction
The classical theory of elliptic integral of the first kind for a Weierstrass family of elliptic curves $X\to S$ (e.g. [Si] Chap.1) defines a period map, which is a multi-valent map from the complement of the discriminant divisor $S\setminus D$ of the base space $S$ of the family to its range domain (so called, period domain). The range is isomorphic to the monodromy covering of $S\setminus D$, i.e. the quotient space of the universal covering $(S\setminus D)^{\sim}$ divided by the kernel of the monodromy representation of $\pi_{1}(S\setminus D,*)$. Then, the Jacobi’s inversion problem for the period map ([ibid] 1.13.) is to ask for a description of the inverse map from the range to the defining domain $S\setminus D$, and its answer were given by Eisenstein series (see [ibid] Section13)
The theory of period integrals for primitive forms [S1] has generalized those classical theory of period integrals to wider classes of geometric families $X\to S$. In particular, the elliptic integrals of Weierstrass type, Jacobi type and Hesse type are explained by the primitive forms associated to the root systems of types $\mathrm{A}_{2},\mathrm{B}_{2}$ and $\mathrm{G}_{2}$, respectively ([S2])). Recently, we find that hyperbolic root systems of rank 2, which do not have geometric origin, admit also the construction of primitive forms and associated period maps ([S3]). We call such period map “virtual” since it does not have corresponding geometric family $X\to S$. Then, in these virtual period maps, we observe some new phenomenon. Namely, an infinite cyclic group, called the winding group, acts on the defining domain and on the monodromy covering space of the period map. It seems to ask for a reformulation of the classical inversion problem.
The purpose of the present note is to illustrate such winding phenomenon by an elementary model of virtual period maps of rank 1. We observe that, in this model, the range is the quotient of the monodromy covering space by the action of the winding group, which we call the winding quotient range. Thus, the classical inversion problem is reformulated in this new setting to a problem to describe the morphism from the winding quotient range to the winding quotient of the defining domain which is actually an elliptic curve for the moduli $q=\exp(2\pi\sqrt{-1}\tau)$.
The new inversion problem asks to construct $q$-multiplicatively periodic functions on the winding quotient range, i.e. it asks to reformulate the classical elliptic function theory from the view point of the winding quotient range. We solve the new problem by introducing $q$-multiplicatively quasi-periodic functions $\mathcal{U}(w,q),\mathcal{V}(w,q)$ and a $q$-multiplicatively periodic function $\mathcal{W}(w,q)$ on the winding quotient range whose pull-backs to the monodromy covering space are the Fourier expansion parts of Weierstrass $\sigma$-function, $\zeta$-function and $\wp$-function, respectively, up to a correction terms studied by Borcherds [B] generated by Eisenstein series $E_{2}$. It seems interesting to note that this answer gives a connection between the classical elliptic function theory with the $q$-shifted factorial function $(w;q)_{\infty}$ in the study of combinatorics and special functions (see Appendix). Two areas of mathematics started from Euler meet again after 3 centuries.
@@
Actually, the pull-back of the function $\mathcal{W}(w,q)$ to the winding covering space has been studied in quite different contexts in mathematical physics either as the propagator $P(z)$ by Dijkgraaf [D] (4.44) on elliptic curves and the propagator $P_{0}^{\infty}(w_{1},w_{2};\tau,\overline{\tau})$ of BCOV theory by Si Li ([L1] Lemma 3.1, [L2] section 4 and [LZ] 3.3), or as the genus-one Bergman kernel function $B(z_{1},z_{2})$ in the topological recursion by Eynard [E] p.339 Examples (c.f. also Tyurin [Ty] p.185 Example. $\omega_{\alpha}$ and Fay [F] (46)). We do not know yet whether such connection of the winding phenomenon with mathmatical physics is common for non-geometric virtual period maps in general or it remains only in these test examples. We need to test further cases [S3, G-S].
@@ 3. Lifting of the map $z^{\tau}$ to the monodromy covering space $\mathbb{C}_{\mathrm{Log}z}$
The naive “inversion problem” in this setting is to give a map from $\mathbb{C}_{w}^{\times}$ to $\mathbb{C}_{z}^{\times}$ making the diagram commutative. However, there does not exists a univalent holomorphic map which makes the diagram commutative, since the analytic continuation of any local inversion map becomes automatically the multi-valued function $z=w^{1/\tau}$ on $\mathbb{C}_{w}^{\times}$. That is, a naive solution of the classical inversion problem for the virtual period map $z^{\tau}$ by a univalent map from the range to the domain does not exist.
@@ 5. Winding of the domain $\mathbb{C}_{z}^{\times}$ of $z^{\tau}$ to the elliptic curve $E_{\tau}$
where the scaling factor $\frac{\tau}{2\pi\sqrt{-1}}$ comes from the scaling of the group $\Gamma_{g}+\Gamma_{w}=\frac{2\pi\sqrt{-1}}{\tau}(\mathbb{Z}+\mathbb{Z}\tau)$ relative to the lattice $\mathbb{Z}+\mathbb{Z}\tau$ for the $\wp$-function. [^3] The realization of $E_{\tau}$ as a plane curve is not unique (e.g. [S2]). The present note has chosen $\mathrm{A}_{2}$-type realization. It would be interesting to study the other cases $\mathrm{B}_{2}$-type and $\mathrm{G}_{2}$-type. Thus, the new inversion problem asks to describe the pair coordinates $(\wp(u,\tau),\wp^{\prime}(u,\tau))$ of the point $e\in E_{\tau}$ in terms of $w=\exp(\tau\mathrm{Log}z)=\exp(2\pi\sqrt{-1}u)$ and $\tau$.
6. $q$-multiplicatively quasi-periodic function $\mathcal{V}(w,q)$ and periodic function $\mathcal{W}(w,q)$
In order to answer to the new inversion problem, we review the classical elliptic function theory from the view point of $q$-periodic, i.e. invariant by the multiplicative action of $q^{\mathbb{Z}}\!:=\!\exp(\tau\Gamma_{g})\!=\!\exp(2\pi\sqrt{\!-1}\tau\mathbb{Z})$, functions on the winding quotient space $\mathbb{C}_{w}\times\mathbb{C}_{q}^{\times}$. We start with $q$-shifted factorial function $(w,q)_{\infty}$, which we regard as the holomorphic function on the winding quotient space. Its logarithmic derivative induces a “positive” $q$-multiplicatively quasi-periodic function $\mathcal{V}_{+}(w,q)$. By making up the “negative” part $\mathcal{V}_{-}(w,q)$ together, we obtain the $q$-quasi-periodic function $\mathcal{V}(w,q)$ of the period 1, which, as a meromorphic function, has simple poles along $1\!-\!wq^{m}\!=\!0$ for $m\!\in\!\mathbb{Z}$. The logarithmic derivative $\mathcal{W}(w,q)$ of $\mathcal{V}(w,q)$ is a $q$-multiplicatively periodic meromorphic function on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, which gives the answer to the new inversion problem. See Appendix for more relationships of them with elliptic function theory.
We start with an infinite product (see Euler [E] p.256) in two variables $w$ and $q$
$(1)\qquad\qquad\qquad\qquad\quad\quad\quad(w;q)_{\infty}\ =\ \prod_{m=0}^{\infty}(1-wq^{m})\qquad\qquad\qquad\qquad\qquad\qquad\qquad\vskip-2.84544pt$
called $q$-shifted factorial function ([Sl] p.88, [GR] p.351). Let us show that the infinite product (1) converges absolutely and compact uniformly to a holomorphic function on the winding quotient space $\mathbb{C}_{w}\times\mathbb{D}_{q}$, whose zero-loci is the union of the divisor $\cup_{n=0}^{\infty}\{1-wq^{n}=0\}$. We shall denote the convergent function again by $(w,q)_{\infty}$.
@@
For any compact set $K\subset\mathbb{C}_{w}\times\mathbb{D}_{q}$, set $c:=\sup\{|w|\mid(w,q)\in K\}$ and $d\!:=\!\sup\{|q|\!\mid\!(w,q)\!\in\!K\}$ so that $\sup\{|wq^{n}|\!\mid\!(w,q)\!\in\!K\}\!\leq\!cd^{n}$. Since $0\!\leq\!d\!<\!1$, the bound is sumable, i.e. $\sum_{n=0}^{\infty}cd^{n}\!=\!c/(1\!-\!d)$. This implies the absolute uniform convergence of (1) to a holomorphic function on $K$ (see [WW] p.30 for infinite products). ∎
As a consequence, we introduce three infinite product functions using $(w,q)_{\infty}$. [^4] Since $\mathcal{U}_{-}(w,q)$ and $\mathcal{U}(w,q)$ (and the following $\mathcal{V}_{-}(w,q)$, $\mathcal{V}(w,q)$ and $\mathcal{W}(w,q)$) have essential singularities at $w=0$, therefore, from now on, we restrict their defining domain to $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$.
$\begin{array}[]{ccccl}\!\qquad\quad\mathcal{U}_{+}(w,q)&:=&(w;q)_{\infty}&\!=\!&\prod_{m=0}^{\infty}(1-wq^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \!(2)\quad\quad\!\mathcal{U}_{-}(w,q)&:=&(w^{-1};q)_{\infty}/(1-w^{-1})&\!=\!&\prod_{m=1}^{\infty}(1-w^{-1}q^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \!\qquad\quad\mathcal{U}(w,q)&:=\!&\!\!\mathcal{U}_{+}(w,q)\ \mathcal{U}_{-}(w,q)&\!=\!&\prod_{m=0}^{\infty}(1\!-\!wq^{m})\prod_{m=1}^{\infty}(1\!-\!w^{-1}q^{m}).\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\end{array}$
Since they converge compact uniformly to holomorphic functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$, we can derivate them by $w$ term-wisely. Thus we obtain three infinite sums
[…]
Note that, by definition directly, we have following formal relations
$(4)\qquad\quad\mathcal{V}_{-}(w,q)=-\mathcal{V}_{+}(w^{-1},q)-\frac{1}{1-w}\quad\text{and}\quad\mathcal{V}(w,q)=\mathcal{V}_{+}(w,q)+\mathcal{V}_{-}(w,q).\qquad\vskip-2.27626pt$
Also by the definition, we obtain following involution property of $\mathcal{U}(w,q)$ and $\mathcal{V}(w,q)$:
$(5)\quad\qquad\qquad\mathcal{U}(w,q)/\mathcal{U}(w^{-1},q)=-w\quad\quad\mathcal{V}(w,q)\ +\ \mathcal{V}(w^{-1},q)\ +\ 1\ =\ 0.\quad\quad\quad\quad\quad$
The absolute compact uniform convergence of (1) implies that these functions are holomorphic except along the singular loci: the divisor $\cup_{n=0}^{\infty}\{1-wq^{n}=0\}$. In order to confirm that they are meromorphic functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$ along the singular loci, we show the compact uniform convergence of the sums (2) in a generalized sense.
Assertion A. (i) The sums (3) converges compact uniformly, in a generalized sense given in the proof, to the meromorphic functions on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$.
(ii) The functions (2) and (3) are $q$-quasi periodic in the following sense:
$\begin{array}[]{cclccc}\quad\qquad\mathcal{U}_{+}(w,q)&=&\mathcal{U}_{+}(qw,q)(1-w)&\mathcal{V}_{+}(w,q)&\!\!\!\!\!\!=\!\!\!\!&\!\!\!\!\!\mathcal{V}_{+}(qw,q)\ \ +\ \ \frac{w}{1-w}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \!\!\!(6)\quad\quad\mathcal{U}_{-}(w,q)&=&\mathcal{U}_{-}(qw,q)(1-w^{-1})^{-1}&\quad\mathcal{V}_{-}(w,q)\quad&\!\!\!\!\!\!=\!\!\!\!&\quad\ \mathcal{V}_{-}(qw,q)\ \ -\ \ \frac{1}{1-w}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\\ \quad\qquad\mathcal{U}(w,q)&=&\ \mathcal{U}(qw,q)(-w)&\mathcal{V}(w,q)&\!\!\!\!\!\!=\!\!\!\!&\!\!\mathcal{V}(qw,q)\ \ \ -\ \ \ \ 1\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\end{array}$
Proof.
(i) We first show only the case $\mathcal{V}_{+}(w,q)$ of (3). That is, in the following proof, (3) means only the case $\mathcal{V}_{+}(w,q)$. First, let us fix following notations.
(a) For $n\in\mathbb{Z}_{\geq 0}$, set $\mathcal{U}_{n}:=\{(w,q)\in\mathbb{C}_{w}\times\mathbb{D}_{q}^{\times}\mid|w|<|q|^{-n-1}\}$.
Then $\mathcal{U}_{n}$ ($n\in\mathbb{Z}_{\geq 0}$) is an increasing sequence such that $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}=\cup_{n\in\mathbb{Z}_{\geq 0}}\mathcal{U}_{n}$.
(b) For any compact subset $K$ of the domain $\mathcal{U}_{n}$ and for any complex valued
continuous function $f$ on $\mathcal{U}_{n}$, we set $f|\!|_{K}:=\mathrm{sup}\{|f(w,q)|\mid(w,q)\in K\}$.
Next, we show that the sum (3)n, obtained from (3) by removing the first $n+1$-terms for $m\!=\!0,1,\cdots,n$, converges uniformly on any compact subset $K$ of the domain $\mathcal{U}_{n}$.
@@
which implies the compact uniform convergence of the sum (3)n.       $\Box$
Thus, (3)n gives a holomorphic function on $\mathcal{U}_{n}$. Then the total sum (3) defines a meromorphic function on $\mathcal{U}_{n}$, having the simple poles along $1\!-\!wq^{m}\!=\!0$ ($m=0,1,\cdots,n$). Recalling that the sequence $\mathcal{U}_{n}$ ($n\in\mathbb{Z}_{\geq 0}$) exhausts $\mathbb{C}_{w}\times\mathbb{D}_{q}^{\times}$, we conclude that the sum (3) defines a meromorphic function on $\mathbb{C}_{w}\times\mathbb{D}_{q}$. We say that (3) converges compact uniformly to the meromorphic function $\mathcal{V}_{+}(w,q)$ in a generalized sense.
The convergence of two other cases $\mathcal{V}_{-}(w,q)$ and $\mathcal{V}(w,q)$ of (3) follows from (4).
Note. By the construction, the principal parts of poles of $\mathcal{V}_{+}(w,q)$ are given by $\frac{1}{1-wq^{m}}$ ($m\in\mathbb{Z}_{\geq 0}$) whose residues are $q^{-m}$.
(ii) The quasi-periodicity (6) of the function $\mathcal{V}_{+}(w,q)$ follows, since the $q$-shift $\mathcal{V}_{+}(wq,q)$ is actually expressed by the sum (3)0 and the 0-th term of (3) is $\frac{w}{1-w}$. The quasi-periodicity of two other cases $\mathcal{V}_{-}(w,q)$ and $\mathcal{V}(w,q)$ follows from (4). ∎
Remark 6.1.
Naive candidates of the functions $\mathcal{U}(w,q)$ and $\mathcal{V}(w,q)$ may be the infinite product $\prod_{m\in\mathbb{Z}}(1-wq^{m})$ and the infinite sum $\sum_{m\in\mathbb{Z}}\frac{wq^{m}}{1-wq^{m}}$, respectively, which are formlly $q$-periodic. However, their negative parts of the index $m\in\mathbb{Z}_{<0}$ do not converge. Therefore, we defined $\mathcal{U}(w,q)$ and $\mathcal{V}(w,q)$ by replacing the negative parts by the convergent $\mathcal{U}_{-}(w,q)=(w^{-1};q)_{\infty}/(1-w^{-1})$ and $\mathcal{V}_{-}(w,q)=\sum_{m\in\mathbb{Z}_{<0}}\frac{1}{1-wq^{m}}$, respectively, as in (2) and (3). These modifications break the (formal) $q$-periodicity of the series, but cause the quasi $q$-periodicity of the functions $\mathcal{U}(w,q)$ and $\mathcal{V}(w,q)$ as in (6). We shall see these two “problems” (i.e. the non-convergency and the non-periodicity) are resolved in the function $\mathcal{W}(w,q)$ introduced in the next paragraph.
We now apply again the action of $w\frac{\partial}{\partial w}$ on $\mathcal{V}(w,q)$. Since the sum (3) is compact uniform convergent (in the generalized sense given in the proof of Assertion A), we can apply the derivation to the series term-wisely so that we obtain a new series:
$\vskip-5.69046pt(7)\qquad\qquad\qquad\qquad\ \ \mathcal{W}(w,q)\ :=\ w\frac{\partial}{\partial w}\mathcal{V}(w,q)\ =\ \sum_{m\in\mathbb{Z}}\frac{wq^{m}}{(1-wq^{m})^{2}}.\qquad\qquad\qquad\ \$
which is compact uniformly convergent for all $m\in\mathbb{Z}$ in the generalized sense. Noting the action $w\frac{\partial}{\partial w}$ is invariant by the $q$-multiplication $w\!\mapsto\!qw$, the $q$-quasi-periodi-city (6) of $\mathcal{V}(w,q)$ implies the $q$-multiplicatively periodicity of the function $\mathcal{W}(w,q)$.
$(8)\qquad\qquad\qquad\quad\ \ \mathcal{W}(qw,q)\ =\ \mathcal{W}(w,q)\qquad\text{and}\qquad\mathcal{W}(w,q)\ =\ \mathcal{W}(w^{-1},q).\quad\qquad\qquad\qquad\qquad\qquad\qquad\qquad$
Assertion B. The function $\mathcal{W}$ on $\mathbb{C}_{w}^{\times}\times\mathbb{D}_{q}^{\times}$ and the $\wp$-function on $\mathbb{C}_{u}\times\mathbb{H}_{\tau}$ are related by the following pull-back relations.
$(9)\qquad\qquad\qquad\qquad\qquad\ \ \wp(u,\tau)\ =\ (2\pi\sqrt{-1})^{2}\ \mathcal{W}(w,q)\ -\ \frac{\pi^{2}}{3}E_{2}(\tau)\quad\quad\quad\quad\qquad\qquad\qquad\quad\vskip-2.84544pt$
where we set
$(10)\qquad\qquad\qquad\qquad\qquad\quad\ w=\exp(2\pi\sqrt{-1}u),\qquad q=\exp(2\pi\sqrt{-1}\tau),\qquad\qquad\qquad\quad\qquad\qquad$
and $E_{2}(\tau)=\frac{1}{\pi^{2}}\sum_{m}\big(\sum_{n}^{\prime}\frac{1}{(m\tau+n)^{2}}\big)$ is the Eisenstein series of weight 2 (here $\sum_{n}^{\prime}$ means n=0n\!=\!0 is omitted if m=0m\!=\!0, see Zagier [Z] p. 19).
@@
For each fixed $\tau\in\mathbb{H}$, the substitution of $w=\exp(2\pi\sqrt{-1}u)$ and $q=\exp(2\pi\sqrt{-1}\tau)$ to $\mathcal{W}(w,q)$ is a doubly periodic meromorphic function in $u\in\mathbb{C}_{u}$ having poles at the lattice $\mathbb{Z}+\mathbb{Z}\tau$, and the principal part of its Laurent expansion at $u=0$ is equal to $\frac{1}{(2\pi\sqrt{-1})^{2}}\frac{1}{u^{2}}$. Then, the difference of the $\wp(u,\tau)$ and the substitution $(2\pi\sqrt{-1})^{2}\mathcal{W}(\exp(2\pi\sqrt{-1}u),\exp(2\pi\sqrt{-1}\tau))$ is a doubly periodic function on $u\in\mathbb{C}_{u}$ without poles so that it is a constant function on $u$. That is, the difference is a holomorphic function depending only on $\tau$.
@@
That is, the pull-back relation holds on the open set $\mathbb{C}_{u}\times\mathbb{H}_{\tau}\setminus\cup_{m\in\mathbb{Z}}\{\Im(u+m\tau)=0\}$. Clearly, this set projects surjectively onto $\mathbb{H}_{\tau}$, implying the equality holds. ∎